OpenStudy (anonymous):

5x + 6 + 6x = 10x + 6 + x

OpenStudy (anonymous):

Take all the X's to one side and the numbers to the other side 5x - 10x + 6x -x = 6 -6 0x = 0 x = 0

OpenStudy (anonymous):

5x + 6 + 6x = 10x + 6 + x 11x+6=10x+6+x 11x+6=11x+6 11x+6-11x=6 0+6=6 6=6

OpenStudy (waleed_imtiaz):

I think that the equation shows that there are infinitely many solutions to the problem... because it shows 0=0 or in some other way 6=6 as u mentioned @amber517 ,

OpenStudy (precal):

11x+6=6+11x solution is all real numbers

OpenStudy (waleed_imtiaz):

@precal Thank you... EM new here....